A silicon pn-junction at T= 300k has a reverse-saturation current of IS=10^16A determine the forward bias diode current for VD=0.55V
1
Expert's answer
2020-06-10T07:49:33-0400

ID = (10-16) [e(0.55/0.026)-1]

= (10-16) [1.53814*109]

= 0.154*10-6

= 0.154uA

Therefore the forward -bias current for VD = 0.55v is ID = 0.154uA